# SoK: Cybersecurity Assessment of Humanoid Ecosystem

### v1

@@ 4.1 Holistic Humanoid Threat Landscape
Defense ID Defense Vector Description Ref. Physical layer - Hardware Integrity and Energy Manipulation P-D1 Secure boot Siddiqui et al. implement hardware root-of-trust for Xilinx FPGAs with authenticated boot bitstreams [90], while NVIDIA Jetson uses on-die BootROM with PKC fuses for signed boot images [60, 59]. This protects humanoid robots using SPI flash boot and FPGA motor controllers from reflashing attacks and privilege escalation. [60, 59, 90] P-D2 Redundant encoders Campeau et al. fused optical encoders with Hall-effect sensors using Kalman filtering on a 2-DoF assistive robot, detecting sensor faults through residual analysis [41]. This applies to humanoids’ balance-critical joints like knees and ankles. [41] P-D3 Rail-level cutouts Ismail et al. described battery-disconnect units with contactor and fast-acting fuses that isolate DC buses during over-current events [32]. For humanoids with 48-800V actuator rails, these cutouts prevent power faults from cascading into compute/control systems. [32] P-D4 Tamper evidence Vidakovic classified conductive break-trace meshes as tamper-evident, providing persistent evidence of tampering that can trigger detection responses. Paley et al. used active PCB meshes measuring trace resistance to detect drilling/cuts and trigger shutdown [65]. In humanoids, this enables tamper evidence and safe-mode responses against physical attacks. [105, 65] P-D5 Leak-resilient mechanics Matlack et al. demonstrated broadband vibration absorption using composite meta-structures, achieving vibration and gear-whine reductions with damping treatments [52]. These retrofittable CLD/meta-material patches apply to humanoid harmonic-drive enclosures to suppress acoustic side-channels. [52] Sensing and Perception Layer - Sensor Fidelity, Privacy, and Signal Integrity SP-D1 Cross-modal verification and temporal coherence You et al. introduced temporal-consistency checks across consecutive LiDAR frames, detecting spoofed objects with >98% accuracy at 41 Hz. Hau et al.’s Shadow-Catcher validates detections by verifying physically consistent 3D shadows, achieving >94% accuracy in 3-21 ms. In humanoids with multiple depth sensors it can prevent spoofed LiDAR data. [116, 29] SP-D2 Coded illumination and pulse signatures Wang et al. introduced pseudorandom-modulation quantum-secured LiDAR, using randomized photon properties and code correlation for ranging against spoofing [109]. Yu et al. developed true-random coded photon-counting LiDAR to generate unique pulse codes [118]. Integrating these into humanoids’ active depth sensors could enable real-time anti-spoofing. [109, 118] SP-D3 Acoustic and EMI hardening Jeong et al. developed a denoising autoencoder that recovered IMU signals under acoustic interference against acoustic injection attack [33]. Kune et al. evaluated shielding, differential inputs, and adaptive filtering to attenuate EMI-borne artifacts [39]. In humanoids they can stabilize IMU-based balance and prevent bias from propagating to gait control. [34, 39] SP-D4 Robust perception via adversarial training Liu et al.’s ARMRO detected  95% of adversarial patches in vision transformers with  1% accuracy loss [44]. Hofman et al.’s X-Detect flags digital/physical patches in real-time with low false positives [31]. These defenses enable humanoids to mask patched regions before visual tracking, SLAM, and object recognition. [44] SP-D5 On-device encryption and data hygiene Kim et al. demonstrated encryption and authentication of low-rate sensor streams [35]. Mi et al. demonstrated accurate face identification despite stripping all information other than required for faceID [55]. In humanoids, these can be used to streams from low rate sensors (IMU, joint states) and preserve privacy. [35, 55] Data Processing Layer - Runtime Pipelines, Memory Safety, and Control Loop Consistency DP-D1 Timing guards Abaza et al. implemented Logical Execution Time (LET) for bounded and scheduled execution of periodic task (e.g., IMU-estimator-controller-actuators) with negligible end-to-end jitter [1]. Ye et al. used LET windows and deadline guards to keep end-to-end timing predictable and prevent late, unstable commands from reaching the actuators [115]. For humanoids, these guards bounded timing and suppressed tardy outputs before actuation, stabilizing balance and gait. [115, 1] DP-D2 Memory safety Schmidt et al. rewrote NAO-class Extended KF/vision in Rust, eliminating common memory-safety bugs for $\approx$6% CPU overhead [83]; Deng et al. evaluated address space layout randomization and stack canarie to contain residual C/C++ faults [19]. For humanoids, porting balance/perception nodes to Rust and hardening legacy nodes reduces crash risk. [83, 19] DP-D3 Race-proof queues Paccagnella et al. exposed deletable audit events in kernel buffers before commit [63]. Ahmad et al.’s HARDLOG sealed events synchronously with tamper-proof devices, eliminating pre-commit races [2]. Wagner et al. used TPM quotes for DDS peer attestation [107]. For humanoids, protected audit devices and attested DDS membership can prevent tampering. [107, 2, 63] DP-D4 Parameter attestation and sanity filters Kuang et al.’s DO-RA attests runtime EKF parameters via MAC checks each cycle, triggering failsafe on mismatch [38]. Tanil et al. use EKF innovation-$\chi^{2}$ gating for sensor exclusion [96], and Bloesch et al. cross-validate exteroceptive pose against IMU/kinematics [11]. This layered approach detects parameter tampering and prevents sensor spoofing. [38, 96, 11] Middleware Layer - Secure Communications, Naming, and Orchestration MW-D1 Cryptographic hardening (DDS-Security / SROS2) SROS2 provided mutual authentication, AES-GCM transport protection, and XML-based authorization policies that reject unauthorized publishers/subscribers on control topics of robot [19]. For humanoids, enforcing SROS2 security features with automated X.509 provisioning constrain remote takeover of motion-critical topics. [19] MW-D2 Intrusion detection and policy enforcement Rivera et al.’s ROSDN enforced policies on each topic policies via SDN firewall on Turtlebot3 [78]. Soriano et al.’s RIPS detected unauthorized subscriptions and commands on TIAGo using rule-based filtering, triggering mitigations limiting video leakage [80]. For humanoids, these policy gates on DDS topics prevents command spoofing and data exfiltration. [78, 80] MW-D3 Information-flow control Pandya et al.’s Picaros enforced decentralized data flow control in ROS 2 using secrecy labels. For humanoids, labeling robot topics confines downstream use and prevents unlabeled nodes from exfiltrating data across distributed deployments. [67]. [67] MW-D4 QoS shaping and freshness enforcement Abaza et al. maintained negligible ROS 2 jitter using LET and reservation servers [1], while DDS Lifespan/Deadline policies bound stale deliveries. Wagner et al. used nonce-fresh attestation for measured peer admission [107]. This combination prevents stale commands from reaching humanoid actuators. [107, 1] Decision-Making Layer - Compromise of Autonomy and Control Logic DM-D1 Redundant and certified decision-making Gandhi et al.’s proposed bounded-time interaction for multi-robot systems, maintaining mission progress despite faulty/malicious robots with modest overhead [26]. Lütjens et al. computed certified action selection under bounded observation perturbations, rejecting actions violating worst-case safety margins at runtime. For humanoids, fleet redundancy and certified action filters prevent compromised modules or noisy sensors from cascading into unsafe behavior. [49]. [49, 26] DM-D2 Adversary-aware planning Shi et al. formulated RMOP that maintains bounded team rewards under $\alpha$ robot dropouts with constant-factor guarantees and online MCTS variants. Soriano et al. built ROS 2 intrusion prevention with topic monitoring and rule-based System Modes mitigations on TIAGo [80]. For humanoids, $\alpha$-tolerant planning and IDS-driven mode switches preserve mission progress and enable safe idle during abnormal communications. [87, 80]
@@
Defense ID Defense Vector Description Ref. DM-D3 Simulated adversary modules (shadow probe) Shi et al. trained a small disturbance policy and ran it alongside the controller; the probe predicted when tiny IMU/force perturbations would make a candidate action unsafe, after which a safety filter discarded the action and logged a reproducible test case [86]. Pinto et al. casted such disturbance agents as adversaries in robust RL [70]. In humanoids, robustness can be improved without retraining by evaluating each step/torque command in parallel and blocking those that failed a worst-case check. [86, 70] DM-D4 Adversarial training and risk-aware policies Xie et al. adversarially trained visual-motor policies with learned visual perturbers for better generalization [112], while Han et al. co-trained with embodied disturbance agents to improve locomotion robustness [23]. Lutjens et al. enforced certified safety filters rejecting unsafe actions under bounded observation noise [49]. For humanoids, combining adversarial training with runtime certified filters hardens control against lighting shifts and sensor errors without retraining. [112, 49, 23] DM-D4 Adversarial training and risk-aware policies Xie et al. adversarially trained visual-motor policies with learned visual perturbers for better generalization [112], while Han et al. co-trained with embodied disturbance agents to improve locomotion robustness [23]. Lutjens et al. enforced certified safety filters rejecting unsafe actions under bounded observation noise [49]. For humanoids, combining adversarial training with runtime certified filters hardens control against lighting shifts and sensor errors without retraining. [112, 49, 23] DM-D5 Poisoning-resilient learning Liu et al. trained behavior-cloning networks with median-of-means loss by splitting demonstrations into groups, averaging each group, then optimizing toward the median average. This down-weights poisoned/low-quality demonstrations while maintaining clean-data performance, directly applicable to humanoids learning from noisy or attackable teleop logs [45]. [45] DM-D6 Backdoor detection Fan proposed PeerGuard that enables agents to cross-check peers and generate explicit reasoning for inconsistencies, detecting poisoned agents with high accuracy in LLM multi-agent evaluations [25]. For humanoids, a peer safety agent could validate action rationales against sensor facts and task intent before execution, flagging compromised policies at runtime. [25] DM-D7 Secure model updates and storage Plappert and Fuchs developed TPM-anchored OTA scheme that authenticated update bundles and derived per-update keys only when TPM policies validated, with all symmetric keys kept inside the TPM. They reported kB-scale metadata overhead with TPM enforcement [71]. For humanoids, signing firmware/ML models, sealing keys to measured state, and using Linux integrity measurement architecture (IMA)/ extended verification module (EVM) or fs-verity prevents forged/replayed packages and blocks unsigned model execution. [71] Application Layer - Task Scripts, APIs, and OTA Operations AP-D1 Signed scripts and runtime monitors Colledanchise et al. generated runtime monitors from behavior tree specifications that detected requirement violations on the IIT R1 robot during service tasks. Linux IMA/EVM or fs-verity enforced BT/YAML integrity with userspace signing, preventing silent edits from loading [16]. For humanoids, this can block tampered task scripts execution and raise faults before unsafe transitions reach actuators. [16] AP-D2 API gateways and intrusion detection ROS 2 enforced per-node permissions with DDS Security’s signed XML (SROS2), restricting topic access rights. Rivera et al. filtered per-topic/GUID flows using an SDN firewall (ROSDN) on Turtlebot3 and enabled policy-based blocking of forged traffic [78]. Soriano et al. built RIPS for ROS 2 with a rule DSL (topic matching, mitigations via System Modes) and demonstrated live prevention on a social robot [80]. For humanoids, strict SROS2 permissions and policy gates on motion-critical topics contain spoofing/exfiltration before affecting balance and gait. [78, 80] AP-D3 Secure OTA and reproducible CI Plappert et al. sealed update-signing keys in TPM with bootloader-enforced signature/version validation, causing only kB-scale metadata overhead [71]. Combined with signed continuous integration (CI) receipts (supply chain levels for software artifacts) and device file checks (fs-verity/IMA), this ensures firmware/ML models match CI builds exactly, preventing forged or replayed packages. This applies to humanoids updating firmware and policies over ROS 2/DDS. [71] AP-D4 Twin hardening and domain randomization Xu et al. showed that limiting per-robot API calls can prevent cloud-planner latency inflation [113], Muratore et al. implemented domain randomization to train policies that tolerate sim-to-real mismatch [57], and Betzer et al.performed runtime checks against a digital twin and triggered a safe stop when live traces violated specified properties [10]. For humanoids, rate-limiting twin APIs to prevent planning stalls, train with domain randomization for sim-to-real robustness, and twin-based runtime verification narrows the reality gap and halts unsafe divergence early. [113, 57, 10] AP-D5 Parameter whitelists and dynamic guards Quarta et al. showed that controller parameter changes on ABB IRB140 that caused accuracy/safety violations without alarms can be prevented by locking safety-critical gains as immutable/signed files (IMA/EVM or fs-verity), restricting parameter access to authorized nodes (SROS2), and adding runtime guards that compare in-memory values to signed baselines with fault trips on divergence [76]. For humanoids, sealing EKF/controller gains and guarding edits blocks instability created from parameter tampering. [76] Social Interface Layer - Speech, Gesture, and the Human-Trust Channel SI-D1 Hardware filtering and signal checks Zhang et al. (EarArray) and Liu et al. (MicGuard) detected inaudible/optical command injections at the microphone front-end before ASR, blocking non-human voice inputs on the human-robot interaction (HRI) channel [123, 47]. For humanoids with always-on voice user interfaces, these SI-layer filters prevent malicious spoken commands from reaching middleware/control, containing cross-layer escalation. [123, 47] SI-D2 Robust ASR and speaker authentication He et al. built an active ultrasonic canceller that transmits guard tones, causing mic non-linearity to fold inaudible commands into audible bands where adaptive filtering detects and subtracts injected content [30]. For humanoids with always-on voice interfaces, mounting head emitters and running cancellers before ASR blocks inaudible triggers and prevents cross-layer escalation.. [30] SI-D3 Anti-spoof vision Liu et al. showed that flashing a short, random light-CAPTCHA and checking the RGB reflections verifies a live face with a standard camera [48]. Wu et al. (DepthFake) harden depth-based liveness by randomizing the projector pattern and enforcing 3D-consistency checks [111]. For humanoids, combining light-CAPTCHA on the RGB path with depth-sensor liveness stops video spoofs at the social-interface layer, preventing credential abuse from propagating to control. [48, 111] SI-D4 Sensor access control and privacy by design Soriano et al.’s RIPS-ROS2 enforced per-topic rules on social-interface endpoints (e.g., /asr/result, /teleop/*) and triggered safe modes, blocking unauthorized commands with $\approx$0.8-1.2s reactions on TIAGo robot [80]. Martin et al. added on-device redaction masking biometric identifiers in HRI streams before off-robot use [51]. For humanoids, pairing RIPS-style SI gates with edge redaction can prevent command abuse and PII leakage beyond the SI layer. [80, 51] SI-D5 Script integrity and trust boundary training Colledanchise et al. generated monitors from signed Behavior-Tree specs that abort illegal transitions in <4ms on IIT R1 robot, stopping tampered task scripts before reaching actuators [16]. Saunderson et al. showed verbal warnings before risky requests cut user compliance, reducing unsafe intent confirmation at voice interfaces [82]. For humanoids, signing BT/YAML with runtime monitoring and requiring SI safety preambles for high-risk intents prevents bad commands from propagating beyond the social interface. [82, 16]

### v2

@@ 4.1 Holistic Humanoid Threat Landscape
Defense ID Defense Vector Description Ref. Physical layer - Hardware Integrity and Energy Manipulation P-D1 Secure boot Siddiqui et al. implement hardware root-of-trust for Xilinx FPGAs with authenticated boot bitstreams [91], while NVIDIA Jetson uses on-die BootROM with PKC fuses for signed boot images [65, 66]. This protects humanoid robots using SPI flash boot and FPGA motor controllers from reflashing attacks and privilege escalation. [65, 66, 91] P-D2 Redundant encoders Campeau et al. fused optical encoders with Hall-effect sensors using Kalman filtering on a 2-DoF assistive robot, detecting sensor faults through residual analysis [13]. This applies to humanoids’ balance-critical joints like knees and ankles. [13] P-D3 Rail-level cutouts Ismail et al. described battery-disconnect units with contactor and fast-acting fuses that isolate DC buses during over-current events [35]. For humanoids with 48-800V actuator rails, these cutouts prevent power faults from cascading into compute/control systems. [35] P-D4 Tamper evidence Vidakovic classified conductive break-trace meshes as tamper-evident, providing persistent evidence of tampering that can trigger detection responses. Paley et al. used active PCB meshes measuring trace resistance to detect drilling/cuts and trigger shutdown [70]. In humanoids, this enables tamper evidence and safe-mode responses against physical attacks. [106, 70] P-D5 Leak-resilient mechanics Matlack et al. demonstrated broadband vibration absorption using composite meta-structures, achieving vibration and gear-whine reductions with damping treatments [55]. These retrofittable CLD/meta-material patches apply to humanoid harmonic-drive enclosures to suppress acoustic side-channels. [55] Sensing and Perception Layer - Sensor Fidelity, Privacy, and Signal Integrity SP-D1 Cross-modal verification and temporal coherence You et al. introduced temporal-consistency checks across consecutive LiDAR frames, detecting spoofed objects with >98% accuracy at 41 Hz. Hau et al.’s Shadow-Catcher validates detections by verifying physically consistent 3D shadows, achieving >94% accuracy in 3-21 ms. In humanoids with multiple depth sensors it can prevent spoofed LiDAR data. [117, 31] SP-D2 Coded illumination and pulse signatures Wang et al. introduced pseudorandom-modulation quantum-secured LiDAR, using randomized photon properties and code correlation for ranging against spoofing [109]. Yu et al. developed true-random coded photon-counting LiDAR to generate unique pulse codes [119]. Integrating these into humanoids’ active depth sensors could enable real-time anti-spoofing. [109, 119] SP-D3 Acoustic and EMI hardening Jeong et al. developed a denoising autoencoder that recovered IMU signals under acoustic interference against acoustic injection attack [36]. Kune et al. evaluated shielding, differential inputs, and adaptive filtering to attenuate EMI-borne artifacts [43]. In humanoids they can stabilize IMU-based balance and prevent bias from propagating to gait control. [37, 43] SP-D4 Robust perception via adversarial training Liu et al.’s ARMRO detected  95% of adversarial patches in vision transformers with  1% accuracy loss [47]. Hofman et al.’s X-Detect flags digital/physical patches in real-time with low false positives [33]. These defenses enable humanoids to mask patched regions before visual tracking, SLAM, and object recognition. [47] SP-D5 On-device encryption and data hygiene Kim et al. demonstrated encryption and authentication of low-rate sensor streams [39]. Mi et al. demonstrated accurate face identification despite stripping all information other than required for face identification [60]. In humanoids, these can be used to streams from low rate sensors (IMU, joint states) and preserve privacy. [39, 60] Data Processing Layer - Runtime Pipelines, Memory Safety, and Control Loop Consistency DP-D1 Timing guards Abaza et al. implemented Logical Execution Time (LET) for bounded and scheduled execution of periodic task (e.g., IMU-estimator-controller-actuators) with negligible end-to-end jitter [1]. Ye et al. used LET windows and deadline guards to keep end-to-end timing predictable and prevent late, unstable commands from reaching the actuators [116]. For humanoids, these guards bounded timing and suppressed tardy outputs before actuation, stabilizing balance and gait. [116, 1] DP-D2 Memory safety Schmidt et al. rewrote NAO-class Extended KF/vision in Rust, eliminating common memory-safety bugs for $\approx$6% CPU overhead [34]; Deng et al. evaluated address space layout randomization and stack canaries to contain residual C/C++ faults [19]. For humanoids, porting balance/perception nodes to Rust and hardening legacy nodes reduces crash risk. [34, 19] DP-D3 Race-proof queues Paccagnella et al. exposed deletable audit events in kernel buffers before commit [68]. Ahmad et al.’s HARDLOG sealed events synchronously with tamper-proof devices, eliminating pre-commit races [2]. Wagner et al. used TPM quotes for DDS peer attestation [107]. For humanoids, protected audit devices and attested DDS membership can prevent tampering. [107, 2, 68] DP-D4 Parameter attestation and sanity filters Kuang et al.’s DO-RA attests runtime EKF parameters via MAC checks each cycle, triggering failsafe on mismatch [42]. Tanil et al. use EKF innovation-$\chi^{2}$ gating for sensor exclusion [100], and Bloesch et al. cross-validate exteroceptive pose against IMU/kinematics [9]. This layered approach detects parameter tampering and prevents sensor spoofing. [42, 100, 9] Middleware Layer - Secure Communications, Naming, and Orchestration MW-D1 Cryptographic hardening (DDS-Security / SROS2) SROS2 provided mutual authentication, AES-GCM transport protection, and XML-based authorization policies that reject unauthorized publishers/subscribers on control topics of robot [19]. For humanoids, enforcing SROS2 security features with automated X.509 provisioning constrain remote takeover of motion-critical topics. [19] MW-D2 Intrusion detection and policy enforcement Rivera et al.’s ROSDN enforced policies on each topic policies via SDN firewall on Turtlebot3 [82]. Soriano et al.’s RIPS detected unauthorized subscriptions and commands on TIAGo using rule-based filtering, triggering mitigations limiting video leakage [94]. For humanoids, these policy gates on DDS topics prevents command spoofing and data exfiltration. [82, 94] MW-D3 Information-flow control Pandya et al.’s Picaros enforced decentralized data flow control in ROS 2 using secrecy labels. For humanoids, labeling robot topics confines downstream use and prevents unlabeled nodes from exfiltrating data across distributed deployments. [72]. [72] MW-D4 QoS shaping and freshness enforcement Abaza et al. maintained negligible ROS 2 jitter using LET and reservation servers [1], while DDS Lifespan/Deadline policies bound stale deliveries. Wagner et al. used nonce-fresh attestation for measured peer admission [107]. This combination prevents stale commands from reaching humanoid actuators. [107, 1] Decision-Making Layer - Compromise of Autonomy and Control Logic DM-D1 Redundant and certified decision-making Gandhi et al.’s proposed bounded-time interaction for multi-robot systems, maintaining mission progress despite faulty/malicious robots with modest overhead [24]. Lutjens et al. computed certified action selection under bounded observation perturbations, rejecting actions violating worst-case safety margins at runtime. For humanoids, fleet redundancy and certified action filters prevent compromised modules or noisy sensors from cascading into unsafe behavior. [51]. [51, 24] DM-D2 Adversary-aware planning Shi et al. formulated RMOP that maintains bounded team rewards under $\alpha$ robot dropouts with constant-factor guarantees and online MCTS variants. Soriano et al. built ROS 2 intrusion prevention with topic monitoring and rule-based System Modes mitigations on TIAGo [94]. For humanoids, $\alpha$-tolerant planning and IDS-driven mode switches preserve mission progress and enable safe idle during abnormal communications. [88, 94]
@@
Defense ID Defense Vector Description Ref. DM-D3 Simulated adversary modules (shadow probe) Shi et al. trained a small disturbance policy and ran it alongside the controller; the probe predicted when tiny IMU/force perturbations would make a candidate action unsafe, after which a safety filter discarded the action and logged a reproducible test case [87]. Pinto et al. cast such disturbance agents as adversaries in robust RL [75]. In humanoids, robustness can be improved without retraining by evaluating each step/torque command in parallel and blocking those that failed a worst-case check. [87, 75] DM-D4 Adversarial training and risk-aware policies Xie et al. adversarially trained visual-motor policies with learned visual perturbers for better generalization [113], while Han et al. co-trained with embodied disturbance agents to improve locomotion robustness [30]. Lutjens et al. enforced certified safety filters rejecting unsafe actions under bounded observation noise [51]. For humanoids, combining adversarial training with runtime certified filters hardens control against lighting shifts and sensor errors without retraining. [113, 51, 30] DM-D5 Poisoning-resilient learning Liu et al. trained behavior-cloning networks with median-of-means loss by splitting demonstrations into groups, averaging each group, then optimizing toward the median average. This down-weights poisoned/low-quality demonstrations while maintaining clean-data performance, directly applicable to humanoids learning from noisy or attackable teleop logs [48]. [48] DM-D6 Backdoor detection Fan proposed PeerGuard that enables agents to cross-check peers and generate explicit reasoning for inconsistencies, detecting poisoned agents with high accuracy in LLM multi-agent evaluations [23]. For humanoids, a peer safety agent could validate action rationales against sensor facts and task intent before execution, flagging compromised policies at runtime. [23] DM-D7 Secure model updates and storage Plappert and Fuchs developed TPM-anchored OTA scheme that authenticated update bundles and derived per-update keys only when TPM policies validated, with all symmetric keys kept inside the TPM. They reported kB-scale metadata overhead with TPM enforcement [76]. For humanoids, signing firmware/ML models, sealing keys to measured state, and using Linux integrity measurement architecture (IMA)/ extended verification module (EVM) or fs-verity prevents forged/replayed packages and blocks unsigned model execution. [76] Application Layer - Task Scripts, APIs, and OTA Operations AP-D1 Signed scripts and runtime monitors Colledanchise et al. generated runtime monitors from behavior tree specifications that detected requirement violations on the IIT R1 robot during service tasks. Linux IMA/EVM or fs-verity enforced BT/YAML integrity with userspace signing, preventing silent edits from loading [17]. For humanoids, this can block tampered task scripts execution and raise faults before unsafe transitions reach actuators. [17] AP-D2 API gateways and intrusion detection ROS 2 enforced per-node permissions with DDS Security’s signed XML (SROS2), restricting topic access rights. Rivera et al. filtered per-topic/GUID flows using an SDN firewall (ROSDN) on Turtlebot3 and enabled policy-based blocking of forged traffic [82]. Soriano et al. built RIPS for ROS 2 with a rule DSL (topic matching, mitigations via System Modes) and demonstrated live prevention on a social robot [94]. For humanoids, strict SROS2 permissions and policy gates on motion-critical topics contain spoofing/exfiltration before affecting balance and gait. [82, 94] AP-D3 Secure OTA and reproducible CI Plappert et al. sealed update-signing keys in TPM with bootloader-enforced signature/version validation, causing only kB-scale metadata overhead [76]. Combined with signed continuous integration (CI) receipts (supply chain levels for software artifacts) and device file checks (fs-verity/IMA), this ensures firmware/ML models match CI builds exactly, preventing forged or replayed packages. This applies to humanoids updating firmware and policies over ROS 2/DDS. [76] AP-D4 Twin hardening and domain randomization Xu et al. showed that limiting per-robot API calls can prevent cloud-planner latency inflation [114], Muratore et al. implemented domain randomization to train policies that tolerate sim-to-real mismatch [Muratore2022], and Betzer et al.performed runtime checks against a digital twin and triggered a safe stop when live traces violated specified properties [8]. For humanoids, rate-limiting twin APIs to prevent planning stalls, train with domain randomization for sim-to-real robustness, and twin-based runtime verification narrows the reality gap and halts unsafe divergence early. [114, Muratore2022, 8] AP-D5 Parameter whitelists and dynamic guards Quarta et al. showed that controller parameter changes on ABB IRB140 that caused accuracy/safety violations without alarms can be prevented by locking safety-critical gains as immutable/signed files (IMA/EVM or fs-verity), restricting parameter access to authorized nodes (SROS2), and adding runtime guards that compare in-memory values to signed baselines with fault trips on divergence [81]. For humanoids, sealing EKF/controller gains and guarding edits blocks instability created from parameter tampering. [81] Social Interface Layer - Speech, Gesture, and the Human-Trust Channel SI-D1 Hardware filtering and signal checks Zhang et al. (EarArray) and Liu et al. (MicGuard) detected inaudible/optical command injections at the microphone front-end before ASR, blocking non-human voice inputs on the human-robot interaction (HRI) channel [122, 49]. For humanoids with always-on voice user interfaces, these SI-layer filters prevent malicious spoken commands from reaching middleware/control, containing cross-layer escalation. [122, 49] SI-D2 Robust ASR and speaker authentication He et al. built an active ultrasonic canceller that transmits guard tones, causing mic non-linearity to fold inaudible commands into audible bands where adaptive filtering detects and subtracts injected content [32]. For humanoids with always-on voice interfaces, mounting head emitters and running cancellers before ASR blocks inaudible triggers and prevents cross-layer escalation.. [32] SI-D3 Anti-spoof vision Liu et al. showed that flashing a short, random light-CAPTCHA and checking the RGB reflections verifies a live face with a standard camera [50]. Wu et al. (DepthFake) harden depth-based liveness by randomizing the projector pattern and enforcing 3D-consistency checks [112]. For humanoids, combining light-CAPTCHA on the RGB path with depth-sensor liveness stops video spoofs at the social-interface layer, preventing credential abuse from propagating to control. [50, 112] SI-D4 Sensor access control and privacy by design Soriano et al.’s RIPS-ROS2 enforced per-topic rules on social-interface endpoints (e.g., /asr/result, /teleop/*) and triggered safe modes, blocking unauthorized commands with $\approx$0.8-1.2s reactions on TIAGo robot [94]. Martin et al. added on-device redaction masking biometric identifiers in HRI streams before off-robot use [54]. For humanoids, pairing RIPS-style SI gates with edge redaction can prevent command abuse and PII leakage beyond the SI layer. [94, 54] SI-D5 Script integrity and trust boundary training Colledanchise et al. generated monitors from signed Behavior-Tree specs that abort illegal transitions in <4ms on IIT R1 robot, stopping tampered task scripts before reaching actuators [17]. Saunderson et al. showed verbal warnings before risky requests cut user compliance, reducing unsafe intent confirmation at voice interfaces [84]. For humanoids, signing BT/YAML with runtime monitoring and requiring SI safety preambles for high-risk intents prevents bad commands from propagating beyond the social interface. [84, 17]